OpenStudy (anonymous):

Medal will be awarded :) workings please - A jogger runs 10 km in 45 minutes, rests for 10 minutes, and then runs another 8 km in 1 hour. The average speed of the jogger over the entire run is A 12.41 km/h B 11.61 km/h C 10.29 km/h D 9 km/h E 9.39 km/h

OpenStudy (anonymous):

i got A :)

OpenStudy (anonymous):

see in order to find the average speed first find the total distance travelled and then divide it by the total time taken

OpenStudy (anonymous):

see total distance =10+0+8=18km total time is 45mins+10mins +1hr =45/60 hr + 10/60 hr+ 1 hr =(45+10+60)/60 hr =115/60 hr hence required average speed = total distance / total time

OpenStudy (anonymous):

check if you get E

OpenStudy (anonymous):

it is e :)

OpenStudy (anonymous):

but if i do 18/1.97 its not exactly e its 9.13?

OpenStudy (anonymous):

115/60 = 1.91666666 =1.92 appx

OpenStudy (anonymous):

ohhhhh epic thanks so much heres your medal :)

OpenStudy (anonymous):

yw but whether u pick me for medal its up to you...(medal isn't important to me..)

OpenStudy (anonymous):

well that is very kind of you i always keep my word :)